OpenStudy (anonymous):

GIVING MEDALS AND WILL BECOME A FAN A circular mirror is surrounded by a square metal frame. The radius of the mirror is 5x. The side length of the metal frame 15x. What is the area of the frame? Write your answer in factored form.

OpenStudy (anonymous):

You just have to multiply the length of the sides, so 5x2 an 15x2 and then add them together to get the radius

OpenStudy (anonymous):

So the radius is 40? @Alimara

OpenStudy (anonymous):

Yes, that would be correct

OpenStudy (anonymous):

Okay thankyou(: @Alimara

OpenStudy (kendricklamar2014):

Area of the Square: \[15x * 15x = 225x^2\]

OpenStudy (kendricklamar2014):

Area of the Mirror: \[3.14(25x)^2\]

OpenStudy (kendricklamar2014):

Area of the Metal: \[225x^2 - 25 \pi x^2 = 25x^2(25-\pi)\]

OpenStudy (kendricklamar2014):

The answer in Factored Form @skyelaaboss is:\[25 x^2 (25 -\pi) \]

OpenStudy (anonymous):

Thankyou(:

OpenStudy (kendricklamar2014):

No Problem